OpenStudy (anonymous):

he sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

n+n+1+n+2=267; 3(n+1)=267 You tell me.

OpenStudy (anonymous):

I don't know..

OpenStudy (anonymous):

I just gave you a solvable formula.

OpenStudy (anonymous):

I don't know how to do it >.<

OpenStudy (anonymous):

3(n+1)=267 Divide by 3 and subtract 1 on both sides.